a second angle of a triangle is twice the smallest angle. The largest an is 20 degrees more than the second angle. Find the three angles of the triangle.

Answer by mananth(16946)

let smallest angle be x
second angle = 2x
largest angle = 2x+20
..
x+2x+2x+20=180
5x+20=180
-20
5x=180-20
5x=160
/5
x= 160/5
x=32 -- smallest angle
..
second angle = 2x = 2*32= 64
third angle = 2x+20 = 2*32+20=84
...
CHECK
32+64+84=180
...
m.ananth@hotmail.ca